# Electromechanical computational model of the human stomach

### v1

@@ Acknowledgements
The authors thank Renate Sachse for valuable discussions on shells and Matthias Mayr for fruitful discussions on the Laplace–Dirichlet problems. They also thank Amadeus Gebauer for the insightful discussions on mesh conversions, constrained mixture theory and prestress. Furthermore, the authors thank Fabiola Rienäcker for valuable discussions on medical image reconstruction. The authors gratefully acknowledge the computing resources provided by the Data Science & Computing Lab at the University of the Bundeswehr Munich. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 350481011, 469698389. AG acknowledges the support from the ERC Consolidator Grant MiGEM #101170592, CUP C83C25000140006.
@@
All the building blocks of the electromechanical finite element model developed and applied in this study are openly available in 4C Multiphysics at https://github.com/4C-multiphysics/4C [32]. Additional software and data that support the findings of this study are available from the corresponding authors upon reasonable request.
@@ A.1 Circumferential and longitudinal collagen and smooth muscle fibers
Collagen and SMCs fiber families ($i\in\{\textup{c},\textup{l}\})$ are modeled as quasi-one-dimensional structures that primarily resist stretch along their respective fiber directions. The strain energy function per unit mass for the $i$-th fiber family is given by the exponential form [40]
@@ A.2 Ground matrix
The remaining structural constituent, the ground matrix ($i=\textup{gm})$, primarily composed of elastin, is modeled as a three-dimensional decoupled isotropic Neo-Hookean material with an isochoric-deviatoric split [33]. The corresponding isotropic passive strain energy function for the ground matrix reads
@@ Appendix B Electrophysiological cell model
Each cell type employs a modified version of the phenomenological two-variable Mitchell-Schaeffer model [82, 83, 16] to characterize ionic exchanges via two state variables: the normalized transmembrane voltage $v^{j}\in[0,1]$ and the recovery (gating) variable $h^{j}\in[0,1]$ for cell type $j\in\{\textup{icc}\,,\ \textup{smc}\}$. Considering the monodomain formulation in Eq. 25, the time evolution of the state variables is governed by the following system of ordinary differential equations
@@
The inward and outward currents are expressed using phenomenological cubic functions [16]
@@ Appendix C Estimation of slow wave conduction velocities
Using elementary trigonometric relations, the activation times of three points can be used to estimate the average conduction velocity and direction, assuming the wavefront is locally planar [68]. This method is illustrated in Figure C.1. For a $30\text{\,}\frac{\mathrm{s}}{}$ window, we extract activation times at all surface nodes of the discretized geometry, defined as the time of the transmembrane potential peak in ICCs. These times are identified directly from the simulation results using the open-source software SciPy [67].
@@
Appendix D Influence of the diffusion coefficient on the propagation pattern
@@
Figure D.1: Sketch of the computational domain of the torus with opening angle $\delta=90^{\circ}$, major radius $R_{1}=$ $159.155\text{\,}\frac{\mathrm{mm}}{}$, minor radius $R_{2}=$ $50.93\text{\,}\frac{\mathrm{mm}}{}$, and wall thickness of $H_{t}=$ $3.5\text{\,}\frac{\mathrm{mm}}{}$. No-flux Neumann boundary conditions are applied at the pyloric sphincter $\Gamma_{\textup{pyl}}$ and pacemaker line $\Gamma_{\textup{prox}}$.
@@
We model the functional distal stomach, excluding the electrically quiescent fundus region, which does not exhibit typical peristaltic contractions [84, 76], using an idealized quarter-torus geometry as shown in Figure D.1. The torus features two openings representing the pyloric sphincter ($\Gamma_{\textup{pyl}}$) and the connection to the proximal stomach ($\Gamma_{\textup{prox}}$). The geometry, generated using Coreform Cubit [59], is dimensioned to mimic anatomical proportions of the human stomach. We choose the mesh sizes based on the convergence study shown in Section 3.2. Consequently, the mesh consists of 373 006373\,006 linear triangular scalar transport elements for the electrophysiology problem.
@@
Figure D.2: Contour plots of spatial distributions of (a) the excitability parameter $a^{\textup{icc}}$, and (b) the ICC diffusion coefficient $\sigma^{\textup{icc}}$.
@@
Figure D.3: Collection of spatial plots illustrating the influence of different spatial distributions of the diffusion coefficient of ICCs and SMCs $\sigma^{j}$ on the propagation of normalized transmembrane potential $v^{\textup{icc}}$. The first row shows the case of a homogeneous distribution, the second row the case of a spatially varying distribution. The colorbar applies to all subplots.
@@
Figure D.4: Estimated conduction velocity $v_{c}$ and propagation patterns of entrained ICC slow waves obtained with the triangulation method. Left: Case with spatially homogeneous diffusion coefficients. Right: Case with spatially heterogeneous diffusion coefficients. In both simulations, excitability is unidirectional. The heterogeneous case produces a spatial gradient in $v_{c}$, enabling wave propagation approximately orthogonal to the longitudinal axis.
@@
We compare electrophysiology simulations using two distributions of the ICC diffusion coefficient $\sigma^{\textup{icc}}$: homogeneous and heterogeneous (as described in Section 2.4.4 and depicted in Figure D.2b). In both cases, a unidirectional excitability parameter distribution $a^{\textup{icc}}$ (see Table E.1 and Figure D.2a) is used to isolate the effects of the diffusion coefficients. Used parameters are listed in Table E.7. Figure D.3 shows the entrained ICC transmembrane potential $v^{\textup{icc}}$ during propagation at selected time points. In the homogeneous case, slow waves do not emerge simultaneously at the pacemaker line despite uniform excitability, instead depolarization occurs first at the pacemaker region. Ring-shaped waves then travel aborally, but the conduction velocity remains nearly constant over the whole domain.
@@
Figure D.4 depicts the resulting differences. Both setups initiate waves at the pacemaker, forming ring-shaped depolarization pattern. However, only the heterogeneous $\sigma^{\textup{icc}}$ distribution produces a conduction velocity gradient sufficient to align wave fronts with geometry curvature. This demonstrates that modulating intrinsic frequencies through the excitability parameter alone is insufficient, and spatially varying diffusion coefficients are required to reproduce physiologically realistic slow wave propagation in curved geometries.
Appendix E Model parameters
Table E.1: Parameters defining the spatial distribution of the excitability $a^{\textup{icc}}$ according to Eq. 31. The functions $f_{\eta}$ for $\eta\in\{l,c\}$ are given by Eq. 32.
@@ Appendix E Model parameters
Table E.2: Baseline electromechanical model parameters used in all simulations unless stated otherwise.
@@
Table E.3: Modified parameter values used in the convergence study. All other parameters are identical to those in Table E.2
@@
Table E.4: Model parameters used for the comparison of finite element formulations.
@@
Table E.5: Initial conditions values used for the electromechanics simulations.
@@
Table E.6: Boundary conditions used for the electromechanics simulations of the realistic stomach. The values correspond to the description in Section 2.5.
@@
Table E.7: Simulation parameter used to study gastric electromechanics.

### v2

@@ Acknowledgements
The authors thank Renate Sachse for valuable discussions on shells and Matthias Mayr for fruitful discussions on the Laplace–Dirichlet problems. They also thank Amadeus Gebauer for the insightful discussions on mesh conversions, constrained mixture theory and prestress. Furthermore, the authors thank Fabiola Rienäcker for valuable discussions on medical image reconstruction. The authors gratefully acknowledge the computing resources provided by the Data Science & Computing Lab at the University of the Bundeswehr Munich. This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 350481011, 469698389. AG acknowledges the ERC Consolidator Grant support from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101170592 — MiGEM.
@@
All the building blocks of the electromechanical finite element model developed and applied in this study are openly available in 4C Multiphysics at https://github.com/4C-multiphysics/4C [4C]. Additional software and data that support the findings of this study are available from the corresponding authors upon reasonable request.
@@ A.1 Circumferential and longitudinal collagen and smooth muscle fibers
Collagen and SMCs fiber families ($i\in\{\textup{c},\textup{l}\})$ are modeled as quasi-one-dimensional structures that primarily resist stretch along their respective fiber directions. The strain energy function per unit mass for the $i$-th fiber family is given by the exponential form [HolzapfelGasserOgden-2012-ConstArterial]
@@ A.2 Ground matrix
The remaining structural constituent, the ground matrix ($i=\textup{gm})$, primarily composed of elastin, is modeled as a three-dimensional decoupled isotropic Neo-Hookean material with an isochoric-deviatoric split [Holzapfel-2001-NonlinearSolidMechanics]. The corresponding isotropic passive strain energy function for the ground matrix reads
@@ Appendix B Electrophysiological cell model
Each cell type employs a modified version of the phenomenological two-variable Mitchell-Schaeffer model [MitchellSchaeffer-2003-TwoCurrentModel, Djabella-2008-CellModel, brandstaeter2018a] to characterize ionic exchanges via two state variables: the normalized transmembrane voltage $v^{j}\in[0,1]$ and the recovery (gating) variable $h^{j}\in[0,1]$ for cell type $j\in\{\textup{icc}\,,\ \textup{smc}\}$. Considering the monodomain formulation in Eq. 25, the time evolution of the state variables is governed by the following system of ordinary differential equations
@@
The inward and outward currents are expressed using phenomenological cubic functions [brandstaeter2018a]
@@ Appendix C Estimation of slow wave conduction velocities
Using elementary trigonometric relations, the activation times of three points can be used to estimate the average conduction velocity and direction, assuming the wavefront is locally planar [Cantwell-2015-Triangulation]. This method is illustrated in Figure C.1. For a $30\text{\,}\frac{\mathrm{s}}{}$ window, we extract activation times at all surface nodes of the discretized geometry, defined as the time of the transmembrane potential peak in ICCs. These times are identified directly from the simulation results using the open-source software SciPy [scipy].
@@
Appendix D Spatial distribution of absolute errors for different element formulations
To complement the maximum error values reported in Table 1, we present the spatial distribution of absolute errors along the longitudinal axis. The absolute error is computed as
$\Delta d_{r}(x)$ $=|d_{r}(x)-d_{r}^{\text{HEX8}}(x)|,$ (56)
$\Delta h_{t}(x)$ $=|h_{t}(x)-h_{t}^{\text{HEX8}}(x)|,$ (57)
where $d_{r}^{\text{HEX8}}$ and $h_{t}^{\text{HEX8}}$ are the reference displacement and thickness profiles obtained from the HEX88 solid formulation.
(a) Radial displacement absolute error $\Delta d_{r}(x)$
(b) Thickness absolute error $\Delta h_{t}(x)$
Figure D.1: Spatial distribution of absolute errors along the longitudinal axis for membrane and 77-parameter shell elements relative to the HEX88 reference solution. Membrane elements show larger errors in regions of high curvature and bending, while shell elements closely follow the reference solution with minor deviations.
The shell formulation closely matches the HEX88 reference, with small deviations mostly occurring near boundaries where locking treatments affect the solution. Membrane elements show significantly larger errors, particularly in regions where out-of-plane bending dominates, consistent with the trends observed in Figures 8b and 8c.
Appendix E Influence of the diffusion coefficient on the propagation pattern
@@
Figure E.1: Sketch of the computational domain of the torus with opening angle $\delta=90^{\circ}$, major radius $R_{1}=$ $159.155\text{\,}\frac{\mathrm{mm}}{}$, minor radius $R_{2}=$ $50.93\text{\,}\frac{\mathrm{mm}}{}$, and wall thickness of $H_{t}=$ $3.5\text{\,}\frac{\mathrm{mm}}{}$. No-flux Neumann boundary conditions are applied at the pyloric sphincter $\Gamma_{\textup{pyl}}$ and pacemaker line $\Gamma_{\textup{prox}}$.
@@
We model the functional distal stomach, excluding the electrically quiescent fundus region, which does not exhibit typical peristaltic contractions [Angeli-2014-ImprovedGutFunction, Athavale-2024-RatGastricSlowWave], using an idealized quarter-torus geometry as shown in Figure E.1. The torus features two openings representing the pyloric sphincter ($\Gamma_{\textup{pyl}}$) and the connection to the proximal stomach ($\Gamma_{\textup{prox}}$). The geometry, generated using Coreform Cubit [coreform2023], is dimensioned to mimic anatomical proportions of the human stomach. We choose the mesh sizes based on the convergence study shown in Section 3.2. Consequently, the mesh consists of 373 006373\,006 linear triangular scalar transport elements for the electrophysiology problem.
@@
Figure E.2: Contour plots of spatial distributions of (a) the excitability parameter $a^{\textup{icc}}$, and (b) the ICC diffusion coefficient $\sigma^{\textup{icc}}$.
@@
Figure E.3: Collection of spatial plots illustrating the influence of different spatial distributions of the diffusion coefficient of ICCs and SMCs $\sigma^{j}$ on the propagation of normalized transmembrane potential $v^{\textup{icc}}$. The first row shows the case of a homogeneous distribution, the second row the case of a spatially varying distribution. The colorbar applies to all subplots.
@@
Figure E.4: Estimated conduction velocity $v_{c}$ and propagation patterns of entrained ICC slow waves obtained with the triangulation method. Left: Case with spatially homogeneous diffusion coefficients. Right: Case with spatially heterogeneous diffusion coefficients. In both simulations, excitability is unidirectional. The heterogeneous case produces a spatial gradient in $v_{c}$, enabling wave propagation approximately orthogonal to the longitudinal axis.
@@
We compare electrophysiology simulations using two distributions of the ICC diffusion coefficient $\sigma^{\textup{icc}}$: homogeneous and heterogeneous (as described in Section 2.4.4 and depicted in Figure E.2b). In both cases, a unidirectional excitability parameter distribution $a^{\textup{icc}}$ (see Table F.1 and Figure E.2a) is used to isolate the effects of the diffusion coefficients. Used parameters are listed in Table F.7. Figure E.3 shows the entrained ICC transmembrane potential $v^{\textup{icc}}$ during propagation at selected time points. In the homogeneous case, slow waves do not emerge simultaneously at the pacemaker line despite uniform excitability, instead depolarization occurs first at the pacemaker region. Ring-shaped waves then travel aborally, but the conduction velocity remains nearly constant over the whole domain.
@@
Figure E.4 depicts the resulting differences. Both setups initiate waves at the pacemaker, forming ring-shaped depolarization pattern. However, only the heterogeneous $\sigma^{\textup{icc}}$ distribution produces a conduction velocity gradient sufficient to align wave fronts with geometry curvature. This demonstrates that modulating intrinsic frequencies through the excitability parameter alone is insufficient, and spatially varying diffusion coefficients are required to reproduce physiologically realistic slow wave propagation in curved geometries.
Appendix F Model parameters
Table F.1: Parameters defining the spatial distribution of the excitability $a^{\textup{icc}}$ according to Eq. 31. The functions $f_{\eta}$ for $\eta\in\{l,c\}$ are given by Eq. 32.
@@ Appendix F Model parameters
Table F.2: Baseline electromechanical model parameters used in all simulations unless stated otherwise.
@@
Table F.3: Modified parameter values used in the convergence study. All other parameters are identical to those in Table F.2
@@
Table F.4: Model parameters used for the comparison of finite element formulations.
@@
Table F.5: Initial conditions values used for the electromechanics simulations.
@@
Table F.6: Boundary conditions used for the electromechanics simulations of the realistic stomach. The values correspond to the description in Section 2.5.
@@
Table F.7: Simulation parameter used to study gastric electromechanics.